### v1

@@ Abstract
[…]
@@ 2.2 Transformer
Yang et al. [23] delved into the CT imaging mechanism and the statistical characteristics of the sinusoidal maps, devising an internal structure loss that encompasses both global and local internal structures to enhance CT image quality. Furthermore, they introduced a sinusoidal map transformation module to more effectively capture sinusoidal map features. By focusing on the internal structure of the sinogram, a feature often overlooked, they significantly reduced image artifacts. Wang et al. [22] introduced a novel locally enhanced window (LeWin) transformer block, utilizing a non-overlapping window-based attention mechanism to reduce the computational demands associated with high-resolution feature maps in capturing local information. Employing this module within a U-net architecture, their approach yielded excellent results in image reconstruction. In another contribution, Wang et al. [21] developed a CT former using an overlapping Window Transformer and Token2Token Dilated strategy within a U-shaped encoder-decoder architecture for LDCT image denoising, achieving remarkable performance. The global perception capabilities of the transformer and the integration of the attention mechanism not only overcome the limitations of convolutional layers’ local feature perception but also enhance global feature extraction and facilitate long-range feature interactions, thereby enriching the information used for LDCT image denoising.
@@ 3.1 Overall architecture
WiTUnet is presented as a U-shaped network architecture, as illustrated in Figure 1 (a). The network consists of encoders, bottleneck, decoders, and an intermediate nested dense block. Specifically, when processing damaged images, namely low-dose computed tomography (LDCT) images represented as $y\in\mathbb{R}^{1\times H\times W}$, WiTUnet initially employs an input embedding layer, a $3\times 3$ convolutional layer with stride and padding set to one, to transform the raw data into feature maps $x\in\mathbb{R}^{C\times H\times W}$. Following the U-shaped architecture [14], these feature maps are processed through $D$ encoder blocks, each comprised of several Window Transformer (WT) blocks, which contain a windowed multi-head self-attention (W-MSA) block. The W-MSA block captures non-local information through its self-attention mechanism and reduces computational complexity using non-overlapping windows, combined with the Local Information Perception Enhancement (LiPe) to precisely capture both local details and global information within the feature maps. Each level of the encoder’s output is down-sampled using a $4\times 4$ convolutional layer with a stride of two as input for the next encoder layer, where the feature maps’ channel count doubles and their dimensions are halved. After $k$ layers of encoding, the feature mapping is represented as $x_{k,0}\in\mathbb{R}^{2^{k}\cdot C\times\frac{H}{2^{k}}\times\frac{W}{2^{k}}}$, where $k\in[0,D)$.
@@ 4.2.3 Quantitative results
Figure 6 facilitates a comparison of the results from six denoising methods: DnCNN, REDCNN, ADNet, NBNet, CTformer, and WiTUnet, illustrating their statistical performance on three key metrics: SSIM, PSNR, and RMSE. DnCNN, as an early denoising network with a simple structure lacking residual learning, exhibits the weakest performance across all metrics, reflecting its limitations in preserving image quality. In contrast, REDCNN, with its integration of an encoder-decoder structure and residual learning, has significantly enhanced its denoising capabilities, particularly evident in the improved SSIM and PSNR scores. Building on this, ADNet optimizes skip connections with convolutional blocks for better feature map alignment, further advancing denoising effectiveness, especially noted in PSNR enhancement. NBNet, with its attention-guided convolutional neural network, strengthens local information capture, manifested in a slight increase in SSIM, although its improvement in RMSE is not as notable as ADNet’s. CTformer introduces a Transformer network, utilizing its non-local information modeling capacity to demonstrate better denoising outcomes across all metrics, yet it marginally lags behind WiTUnet in RMSE, suggesting room for improvement in local detail restoration. WiTUnet, combining the global modeling strength of window Transformers with local information capture, and augmented by nested dense block alignment, shows the best performance across all evaluation metrics, particularly in structural preservation and error minimization. These results not only showcase the formidable capabilities of advanced denoising network architectures but also reflect the importance of architectural design in influencing denoising quality.
@@ 4.2.4 Ablation study
The ablation study was designed to validate the efficacy of the proposed LiPe module and the nested dense block in enhancing the denoising performance of the network. The impact of each module on denoising was assessed using SSIM, PSNR, and RMSE as performance metrics. Here, symbol ① represents the LiPe module, while ② denotes the nested dense block. The absence of ① implies the utilization of an MLP as the FFN, and the absence of ② indicates the use of traditional skip connections as found in a conventional Unet architecture. From our observations, it was found that the network’s performance was at its lowest across all metrics when neither ① nor ② was employed. Introducing ① resulted in a noticeable improvement in all evaluative metrics, suggesting an enhanced focus on local information processing by the network. Similarly, the incorporation of ② led to a higher degree of feature map alignment between the encoder and decoder, thus enabling the network to more effectively reconstruct the FDCT image from the LDCT data. The combined use of both ① and ② yielded the highest results across SSIM, PSNR, and RMSE, indicating that the integration of these two modules not only improved the network’s local feature extraction capabilities but also facilitated better alignment of information between the encoder and decoder feature maps.‘
@@ 5 Discussions
The nested dense blocks within the WiTUnet architecture play a critical role in enhancing the alignment of feature maps between the encoder and the decoder. This innovative design element is instrumental in refining the network’s capability to seamlessly integrate and synchronize the information flow across these crucial components of the denoising process. By facilitating more precise alignment, the nested dense blocks ensure that the encoder and decoder work in harmony, which is essential for maintaining the integrity and continuity of the image features. This alignment is vital for recovering high-fidelity details from LDCT images, enabling the network to effectively suppress noise while preserving essential diagnostic information. This sophisticated architectural innovation not only improves the quality of the denoising output but also significantly contributes to the overall accuracy and effectiveness of medical image analysis facilitated by WiTUnet.
@@ 6 Conclusion
[…]

### v2

@@ Abstract
Low-dose computed tomography (LDCT), Denoising, Convolutional Neural Network, Transformer
@@ 2.2 Transformer
Yang et al. [20] delved into the CT imaging mechanism and the statistical characteristics of the sinusoidal maps, devising an internal structure loss that encompasses both global and local internal structures to enhance CT image quality. Furthermore, they introduced a sinusoidal map transformation module to more effectively capture sinusoidal map features. By focusing on the internal structure of the sinogram, a feature often overlooked, they significantly reduced image artifacts.
Wang et al. [8] introduced a novel locally enhanced window (LeWin) transformer block, utilizing a non-overlapping window-based attention mechanism to reduce the computational demands associated with high-resolution feature maps in capturing local information. Employing this module within a U-net architecture, their approach yielded excellent results in image reconstruction. In another contribution, Wang et al. [7] developed a CT former using an overlapping Window Transformer and Token2Token Dilated strategy within a U-shaped encoder-decoder architecture for LDCT image denoising, achieving remarkable performance. The global perception capabilities of the transformer and the integration of the attention mechanism not only overcome the limitations of convolutional layers’ local feature perception but also enhance global feature extraction and facilitate long-range feature interactions, thereby enriching the information used for LDCT image denoising.
@@ 3.1 Overall architecture
WiTUnet is presented as a U-shaped network architecture, as illustrated in Figure 1 (a). The network consists of encoders, bottleneck, decoders, and an intermediate nested dense block. Specifically, when processing damaged images, namely low-dose computed tomography (LDCT) images represented as $y\in\mathbb{R}^{1\times H\times W}$, WiTUnet initially employs an input embedding layer, a $3\times 3$ convolutional layer with stride and padding set to one, to transform the raw data into feature maps $x\in\mathbb{R}^{C\times H\times W}$. Following the U-shaped architecture [11], these feature maps are processed through $D$ encoder blocks, each comprised of several Window Transformer (WT) blocks, which contain a windowed multi-head self-attention (W-MSA) block. The W-MSA block captures non-local information through its self-attention mechanism and reduces computational complexity using non-overlapping windows, combined with the Local Information Perception Enhancement (LiPe) to precisely capture both local details and global information within the feature maps. Each level of the encoder’s output is down-sampled using a $4\times 4$ convolutional layer with a stride of two as input for the next encoder layer, where the feature maps? channel count doubles and their dimensions are halved. After $k$ layers of encoding, the feature mapping is represented as $x_{k,0}\in\mathbb{R}^{2^{k}\cdot C\times\frac{H}{2^{k}}\times\frac{W}{2^{k}}}$, where $k\in[0,D)$.
@@ 4.2.3 Quantitative results
Figure 6 facilitates a comparison of the results from six denoising methods: DnCNN, REDCNN, ADNet, NBNet, CTformer, and WiTUnet, illustrating their statistical performance on three key metrics: SSIM, PSNR, and RMSE. DnCNN, as an early denoising network with a simple structure lacking residual learning, exhibits the weakest performance across all metrics, reflecting its limitations in preserving image quality. In contrast, REDCNN, with its integration of an encoder-decoder structure and residual learning, has significantly enhanced its denoising capabilities, particularly evident in the improved SSIM and PSNR scores. Building on this, ADNet optimizes skip connections with convolutional blocks for better feature map alignment, further advancing denoising effectiveness, especially noted in PSNR enhancement. NBNet, with its attention-guided convolutional neural network, strengthens local information capture, manifested in a slight increase in SSIM, although its improvement in RMSE is not as notable as ADNet?s. CTformer introduces a Transformer network, utilizing its non-local information modeling capacity to demonstrate better denoising outcomes across all metrics, yet it marginally lags behind WiTUnet in RMSE, suggesting room for improvement in local detail restoration. WiTUnet, combining the global modeling strength of window Transformers with local information capture, and augmented by nested dense block alignment, shows the best performance across all evaluation metrics, particularly in structural preservation and error minimization. These results not only showcase the formidable capabilities of advanced denoising network architectures but also reflect the importance of architectural design in influencing denoising quality.
@@ 4.2.4 Ablation study
The ablation study was designed to validate the efficacy of the proposed LiPe module and the nested dense block in enhancing the denoising performance of the network. The impact of each module on denoising was assessed using SSIM, PSNR, and RMSE as performance metrics. Here, symbol ① represents the LiPe module, while ② denotes the nested dense block. The absence of ① implies the utilization of an MLP as the FFN, and the absence of ② indicates the use of traditional skip connections as found in a conventional Unet architecture. From our observations, it was found that the network’s performance was at its lowest across all metrics when neither ① nor ② was employed. Introducing ① resulted in a noticeable improvement in all evaluative metrics, suggesting an enhanced focus on local information processing by the network. Similarly, the incorporation of ② led to a higher degree of feature map alignment between the encoder and decoder, thus enabling the network to more effectively reconstruct the FDCT image from the LDCT data. The combined use of both ① and ② yielded the highest results across SSIM, PSNR, and RMSE, indicating that the integration of these two modules not only improved the network’s local feature extraction capabilities but also facilitated better alignment of information between the encoder and decoder feature maps.
@@ 5 Discussions
The nested dense blocks within the WiTUnet architecture play a critical role in enhancing the alignment of feature maps between the encoder and the decoder. This innovative design element is instrumental in refining the network?s capability to seamlessly integrate and synchronize the information flow across these crucial components of the denoising process. By facilitating more precise alignment, the nested dense blocks ensure that the encoder and decoder work in harmony, which is essential for maintaining the integrity and continuity of the image features. This alignment is vital for recovering high-fidelity details from LDCT images, enabling the network to effectively suppress noise while preserving essential diagnostic information. This sophisticated architectural innovation not only improves the quality of the denoising output but also significantly contributes to the overall accuracy and effectiveness of medical image analysis facilitated by WiTUnet.
@@ 6 Conclusion
WiTUnet introduces an efficient denoising network structure for LDCT images, distinguished by its innovative nested dense blocks and redesigned skip pathway connections. This architecture not only improves the flow of information between the encoder and decoder but also aligns the semantic levels of the feature maps across different stages, offering the optimizer an advantage in solving optimization challenges. Notably, WiTUnet employs nested skip pathway connections and dense convolutional blocks, considerably enhancing the network’s capacity to capture and integrate both global and local features. During the denoising of LDCT images, this structure adeptly balances detail restoration within feature maps with noise suppression, which is of significant importance for increasing the accuracy of subsequent clinical diagnoses. Moreover, with the incorporation of LiPe and W-MSA, WiTUnet further improves its sensitivity to details and global comprehension of complex image content. In summary, WiTUnet demonstrates cutting-edge potential and clinical applicability in the domain of LDCT image processing.